OpenStudy (anonymous):

A student scored 83 and 91 on her first two quizzes. Write and solve a compound inequality to find the possible values for a third quiz score that would average between 85 and 90. Inclusive. A. 85 <= 83 + 91 + n (over) 3<= 90;81<=n<=96 B. 85 <= 83 + 91 (over) 2 <= n; -2 <= n <= 3 C. 90 <= 83 + 91 + n (over) 3 <=85;96<=n<=81 D. 83 <= 83 + 91 + n (over) 3 <=90;73<=n<=94

OpenStudy (campbell_st):

well the lower limit of the mean is 85 and upper mean is 90 so \[85 \le \frac{83 + 91 + n}{3} \le 90\] so just solve for n...

OpenStudy (anonymous):

i think it is A because it cant be less than the average cant be less than 85 but it can be greater than 90 so D is automatically out because it wants the average at an 83 and B is also out because it says that the average is between 85 and 3 so that narrows it down to 2 answer choices.. hope i helped. (:

OpenStudy (anonymous):

Thank you both :)

OpenStudy (anonymous):

Np :P